OpenStudy (anonymous):

Write the equation of a circle with a center at (1, 2) and a diameter of 15.

OpenStudy (netlopes1):

\[C:(x-1)^{2}+(y-2)^2=( \dfrac{15}{2})^{2}\]

OpenStudy (netlopes1):

Bacause the general formla is \[(x-q)^2 + (y-p)^2 = r^2\], where C(q,p) and radio is r, ok?

OpenStudy (netlopes1):

Sorry, replace "radio" to "radius"....

OpenStudy (jango_in_dtown):

diameter=15 so radius is 15/2 equation of the circle radius r and centre (a,b) is given by (x-a)^2+(y-b)^2=r^2 So we have here cente (1,2) and radius 15/2 so the equation is (x-1)^2+(y-2)^2=(15/2)^2